1. Suppose a company has two mutually exclusive projects, both of which are three years in length. Project A has an initial outlay of $7,000 and has expected cash flows of $3,000 in year 1, $4,000 in year 2, and $4,000 in year 3. Project B has an initial outlay of $10,000 and has expected cash flows of $2,000 in year 1, $4,000 in year 2, and $5,000 in year 3. The required rate of return is 12% for projects at this company. What is the net present value for the best project? (Answer to the nearest dollar.)

2. Suppose a company has proposed a new 5-year project. The project has an initial outlay of $178,000 and has expected cash flows of $32,000 in year 1, $40,000 in year 2, $57,000 in year 3, $68,000 in year 4, and $74,000 in year 5. The required rate of return is 14% for projects at this company. What is the profitability index for this project? (Answer to the nearest hundredth, e.g. 1.23)

3. Suppose a company has proposed a new 4-year project. The project has an initial outlay of $17,000 and has expected cash flows of $8,000 in year 1, $9,000 in year 2, $10,000 in year 3, and $14,000 in year 4. The required rate of return is 15% for projects at this company. What is the profitability index for this project? (Answer to the nearest hundredth, e.g. 1.23)
Answer #1

1.

NPV of Project A=-7000+3000/1.12+4000/1.12^2+4000/1.12^3=1714.46793
NPV of Project B=-10000+2000/1.12+4000/1.12^2+5000/1.12^3=-1466.608965

NPV of the best project=1714.46793

2.
=1+(-178000+32000/1.14+40000/1.14^2+57000/1.14^3+68000/1.14^4+74000/1.14^5)/178000=0.988859511

3.
=1+(-17000+8000/1.15+9000/1.15^2+10000/1.15^3+14000/1.15^4)/17000=1.667148383